OpenStudy (anonymous):

Simplify square root 72 minus square root 27 plus 2 times the square root of 8. negative 30 times the square root of 6 7 square root of 5 negative 2 times the square root of 3 plus 5 times the square root of 2 10 times the square root of 2 minus 3 times the square root of 3

OpenStudy (anonymous):

\[\sqrt{72}-\sqrt{72}+2\sqrt{8}\]

OpenStudy (anonymous):

The first answer choice is \[-30\sqrt{6}\]

OpenStudy (anonymous):

The second is \[7\sqrt{5}\]

OpenStudy (anonymous):

\[-2\sqrt{3}+5\sqrt{2}\]

OpenStudy (anonymous):

the last one is \[10\sqrt{2}-3\sqrt{3}\]